EVA Report – February 23rd

Crew 263 EVA Report 23-02-2022

EVA # 2

Author: Cerise Cuny

Purpose of EVA: battery change for the atmospheric experiment.

Start time: 9:00 am
End time: 9:37 am

Narrative: Because of weather conditions, we only went out to change the batteries of the atmospheric experiments. We checked the rovers’ batteries, just in case. Everything went smoothly, no one slipped.

Destination: Atmospheric experience site (at the intersection between Cow Dung Road and Entrance Road).

Coordinates (use UTM NAD27 CONUS): 518288 E, 4250586 N

Participants: Cerise Cuny, Léa Rouverand, Valentine Bourgeois, Mathéo Fouchet

Road(s) and routes per MDRS Map: Entrance Road.

Mode of travel: Foot

Research Report – February 23rd

Name: Marine Prunier
Crew: 263
Date: 23 Feb 2022

Robotic Telescope Requested (choose one): MLC-RCOS16
Objects to be Imaged this Evening: M81, Horsehead Nebula
Images submitted with this report: None
Problems Encountered: The observation of TOI 1555.01 was canceled in Skynet.


Solar Features Observed: n/a
Images submitted with this report: n/a
Problems Encountered: n/a

Journalist Report – February 23rd

Sol 2 – A day on Mars’ North pole

Author: Nicolas Wattelle

Today was supposed to be a very busy day, and even if it didn’t go as planned, I will be happy to sleep well tonight.

The EVA program was completely full for the whole morning. Changing batteries for the field mill, putting back the LOAC that we took out yesterday (because it cannot stand too wet weather conditions…), installing the antenna for the HF waves propagation study, first look at Candor Chiasma in order to find a great spot to begin the geological experiment (and use LIBS for the first time!).

But it felt like we moved to Mars’ north pole during the night: when we woke up, there was snow everywhere, about 4cm around the MDRS. At first sight, this was really beautiful to see, but we changed our minds when we understood it was compromising our plans… Doing an EVA as long as we had planned would have been risky, so it was decided that the EVA would just aim to change the battery of the field mill (this was the only thing necessary, waiting another day to change the battery would have put at risk its lifetime…).

Also, the snow completely covered our solar panels. Usually, these work in pair with a generator, and they provide enough energy to endure a day at the MDRS, with a crew living and working. So, we had to cut our electrical consumption for the day, keeping only essential devices plugged. This explains the pictures of preparing for the EVA, only using the light coming through the narrow windows to see.

I think this is what trying to live on Mars would look like, everyone will have to get out of their comfort zone, living day to day with the recurrent hazards… But Crew 263 is here to live this experiment, so we accept our fate and hope for better days!

During the afternoon, we were almost all together in the upper deck of the Hab. We made rotations around different experiments, in which sometimes we are operators and sometimes, subjects. On my side, I was mainly focused on producing the first podcast episode, for curious students (and all other curious!) we collaborate with. We hope people on Earth are interested in our mission.

We tried for the first time the ultrasound experiment using a CNES protocol and a Sonoscanner device. It was really interesting, the method is very smart: with one hand, we handle the ultrasound probe, on the other a smartphone with Augmented Reality software. It shows where to place the probe and is thus like a game: we have to put 4 spheres in 4 cubes, and when it is done, we record the sequence. None of us are doctors, but thanks to this protocol we can manage to get great pictures of our organs! We are really grateful to CNES for letting us participate in the conception of this futuristic device!

On another topic, I have to confess something. We are not only 7 in this station… Since we have arrived, someone is there in the Hab upper deck, looking at us all day. He proudly stands at the window. His name is Tomarsto, and he is always there, during great moments and hard times. We hope he will stay there until the end of the simulation, and even more! Take a look at him in the picture of the day!

Operations Report – February 23rd

Crew 263 Operations Report 23-02-2022

SOL: 2

Name of person filing report: Mathéo Fouchet

Non-nominal systems: NA

Notes on non-nominal systems: NA


Spirit rover used: No
Hours: 157.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: handled by Atila

Opportunity rover used: No
Hours: 84.4
Beginning charge: 100%
Ending charge: 100%
Currently charging: handled by Atila

Curiosity rover used: No
Hours: 168.4Beginning charge: 100%

Ending charge: 100%
Currently charging: handled by Atila

Perseverance rover used: NoHours: 228.1
Beginning charge: 100%
Ending charge: 100%
Currently charging: handled by Atila

General notes and comments: NA

Summary of Hab operations:

WATER USE: 17 gallons
Water (static tank): 458 gallons
Water (loft tank): 38 gallons
Water Meter: 01560796 units

Static to Loft Pump used: Yes
Static tank pipe heater: On
Static tank heater: On

Toilet tank emptied: Yes

Summary of internet: NA

Summary of suits and radios: I have changed the fuse of suit #10 so that it can work correctly thanks to the advice of Scott Davis

Summary of GreenHab operations:

WATER USE: 4 gallons at 7:30 am and 4 gallons at 4:45 pm
Heater: On
Supplemental light: Off (will turn on automatically at 7:30 pm)
Harvest: NA

Summary of ScienceDome operations:

Dual split: Off
SOC: 86% at 6pm

Summary of RAM operations: NA

Summary of any observatory issues: NA

Summary of health and safety issues: See HSO report

Questions, concerns and requests to Mission Support: NA

Sol Summary – February 23rd

Crew 263 Sol Summary Report 23-02-2022

Sol: 2

Summary Title: Tiresome cognitive experiments.

Author’s name: Cerise Cuny

Mission Status: Nominal

Sol Activity Summary: This morning we woke up with a Mars polar landscape, and immediately knew that the EVA would be jeopardized because of the snow. In the end, it was decided with mission support that we would only get out to change the batteries of the atmospheric experiment. It was a brief EVA (40 minutes); the landscape was beautiful. The Crew members that stayed in the Hab performed the ultrasound experiment for the CNES. After the EVA, our Crew Engineer repaired the faulty suit and swept the corridors.

In the afternoon, we went through our experiments just as planned. Mathéo and Marine did our human factors experiment, we all were subjects and operators for the ultrasound experiment, and Nicolas and I were subjects of a cognitive experiment. This last one was particularly exhausting: we had to teleoperate a rover lying on the side (among other body positions). I was totally disoriented, I flipped the rover, and it took me 15 minutes to get it back on its wheels! Nicolas is also frustrated because he flipped the rover too: this experiment may have an impact on our mood.

Look Ahead Plan: If the weather allows it, we will set up the antenna during the EVA to be able to start communicating with Toulouse, France (simulating another mars station) and explore a predefined geological area of interest for LIBS analysis. The afternoon will be less intense than today, just a few experiments and hopefully some rest.

Anomalies in work: None to report.

Weather: Cloudy, windy.

Crew Physical Status: Good.

EVA: One performed, everything went fine.

Reports to be filed: Sol Summary, 2 EVA Requests, EVA Report, Operations Report, Journalist Report, Astronomy Report, HSO Report.

Support Requested: We are out of garbage bags, and digesters. Just in case it did not reach the right person yesterday: it would be more convenient for us to have a second working GPS for tomorrow’s EVA as we need to record the GPS information of the observed point. Currently, only one GPS is working, it would be better to have another one for redundancy.

Copyright © The Mars Society. All rights reserved. | Main Site